# QCD Equation of State at very high temperature: computational strategy, simulations and data analysis

### v1

@@ Abstract
We present a detailed account of the theoretical progress and the computational strategy that led to the non-perturbative determination of the QCD Equation of State (EoS) at temperatures up to the electroweak scale reported in [Phys. Rev. Lett. 134, 201904 (2025)]. The two key ingredients that make such a calculation feasible with controlled uncertainties are: (i) the definition of lines of constant physics through the running of a non-perturbatively defined finite-volume coupling across a wide range of energy scales, and (ii) the use of shifted boundary conditions which allow a direct determination of the entropy density thus without the need for a zero-temperature subtraction. Considering the case of QCD with $N_{f}=3$ massless flavours in the temperature interval between 3 GeV and 165 GeV, we describe the numerical strategy based on integrating in the bare coupling and quark mass, the perturbative improvement of lattice observables, the optimization of numerical simulations, and the continuum extrapolation. Extensive consistency checks, including finite-volume and topological-freezing effects, confirm the robustness of the method. The final results have a relative accuracy of about $1\%$ or better, and the errors are dominated by the statistical fluctuations of the Monte Carlo ensembles. We also compare our non-perturbative results with predictions from standard and hard thermal loop perturbation theory showing that at the level of $\%$-precision contributions beyond those known, including non-perturbative ones due to ultrasoft modes, are relevant up to the highest temperatures explored. The methodological framework is general and readily applicable to QCD with four and five massive quark flavours and to other thermal observables, paving the way for systematic non-perturbative studies of thermal QCD at very high temperatures.
@@ I Introduction
The study of the thermodynamics of Quantum Chromodynamics (QCD) is central to our understanding of strongly-interacting matter under extreme conditions, with direct implications for cosmology [1, 2] and heavy-ion physics [3, 4]. A key challenge is the reliable determination of equilibrium properties from first principles across energy scales where perturbation theory is either ineffective or insufficiently accurate. Lattice QCD has proven to be the most powerful tool in this domain with full control on the statistical and on the systematic effects. However, traditional numerical strategies for computing the Equation of State (EoS) by Monte Carlo simulations face severe limitations when extended to temperatures higher than 1 GeV or so. In particular, ultraviolet subtractions and uncertainties in the setting of the bare parameters to determine the lines of constant physics, complicate the realization of this computation and the extraction of results holding in the continuum limit.
To overcome these obstacles, we have developed novel methodologies. They combine non-perturbative renormalization techniques based on the running of a renormalized finite-volume gauge coupling up to very high energies with shifted boundary conditions in the temporal direction, allowing to directy determinate the entropy density which does not need zero-temperature subtractions. This framework allows precise simulations at high temperatures while keeping the computational cost under control and providing a simple and robust method to extrapolate the results to the continuum limit at a given physical temperature $T$. The approach is general and can be systematically applied to QCD with an arbitrary number of quark flavours, including heavy ones, without further theoretical adjustments.
@@ V.1 Finite volume effects
are defined as in Section 3 of Ref. [10]. Notice that the lattice determination of the matrix elements ${\langle T_{\mu\nu}^{G,\{6\}}\rangle}_{{\boldsymbol{\xi}}}$ and ${\langle T_{\mu\nu}^{F,\{6\}}\rangle}_{{\boldsymbol{\xi}}}$ is significantly less demanding compared to the determination of the entropy density using Eq. (6).
@@ VI Equation of State
The main non-perturbative results are the values of the normalized entropy density $s/T^{3}$ of continuum QCD reported in Table 5 with the fit label id3, corresponding to the 9 temperatures from 3 GeV to 165 GeV of Table 1. These values are shown as black dots in Figure 5 as a function of the strong coupling squared $\hat{g}^{2}$ defined above. These results can be fitted to obtain an analytic espression for the temperature dependence of the entropy density, and thus for the other thermodynamic functions in the EoS, namely the pressure and the energy density.
@@ VI.1 Temperature dependence of the entropy density
[…]
@@
Figure 7: Left: Normalized entropy density $s/T^{3}$ in the continuum limit as a function of $\hat{g}^{2}$. The blue band is the best parametrization for $T\geq 500$ MeV, and includes the point at $T=500$ MeV obtained from Refs. [49, 50, 32] (open circle marker). Right: EoS as a function of the temperature foro $T\geq 500$ MeV.
@@ VII Conclusions
A detailed analysis of discretization effects, of finite-volume corrections and of topological contributions confirms the robustness of our results. We have also performed a comparison with perturbative predictions, including both standard and hard thermal loop resummations. Our findings show that, at our level of accuracy, higher-order contributions in $\hat{g}^{2}$ beyond those known, including non-perturbative ones due to ultrasoft modes, are relevant even at temperatures as high as 165 GeV. Furthermore, our results are consistent with existing lattice data at lower temperatures, down to 500 MeV, and a smooth interpolation can be drawn across the entire range.
The methodology discussed in this paper is broadly applicable and can be extended to QCD with four or five massive quark flavours. By disentangling the methodological framework from its first physical application, this work complements earlier results and sets the stage for a systematic program of non-perturbative studies of QCD in the high-temperature regime, with potential impact on cosmology, heavy-ion phenomenology, and the theoretical development of thermal field theory.
@@ Appendix B Free-energy density in lattice perturbation theory
In this Appendix we discuss the results of the free-energy density at one-loop order in lattice perturbation theory, evaluated in the infinite spatial volume limit. These expressions, taken from Ref. [10] (in particular Appendices E and F) are used for the perturbative improvement of the entropy density computed non-perturbatively on the lattice.
@@ B.3.2 Integrals at one-loop order
In left panel of Figure 8 we compare the approach to the thermodynamics limit of $K_{1}^{\prime}$ (orange square markers) and $K_{1}$ (blue circle markers). For all practical purposes we can quote as the infinite volume limit the value of $K_{1}^{\prime}$ at the largest volume, whose discrepancy with respect to the extrapolation is well below the permille level.
The leading finite size corrections of the integral $K_{2}$ are of order $O(1/L^{3})$ and they are negligible at the largest volume $L/a=288$ that we have considered. Thus, we take that value as its thermodynamic limit.
@@ B.4 Infinite volume limit of fermionic contributions
We extrapolate to the thermodynamic limit the tree-level fermionic free-energy density $f^{F(0)}$, Eq. (58), and the one-loop order mass counterterm, Eq. (67), by increasing the system size in the spatial directions from $L/a=96$ up to $L/a=288$ at fixed $L_{0}/a$. The convergence to the infinite volume limit is very fast because, at asymptotically high temperatures, fermions develop a thermal mass $M_{\rm gap}\sim\pi T$ and finite volume effects are exponentially suppressed as $e^{-M_{\rm gap}L}$ [9]. At the largest volume, finite size effects are well below the permille level, and in practice the results can be considered as the values in the thermodynamic limit.
@@
whose integrand has a pole when the gluonic momentum vanishes $q\to 0$ and $p\sim k$. Following the same strategy as for the computation of $K_{1}$ in the previous Section, we introduce the subtracted integral
@@ C.2 Relative error
In the presence of shifted boundary conditions, the temperature of the thermal system depends on the shift through the relation $T^{-1}=L_{0}\sqrt{1+{{\boldsymbol{\xi}}}^{2}}$. At a given $L_{0}$, the temperature changes by less than a factor 2 with respect to the reference case ${{\boldsymbol{\xi}}}=(1,0,0)$ when the shift is varied within the values listed in Eq. (78). Since the entropy density increases slowly from 3 GeV to 165 GeV (about $3\%$, see Table 5), at these temperatures and at a given $L_{0}$ we can assume $s/T^{3}$ to be approximately constant when the shift is changed among the values under investigation. Using this assumption in Eq. (77), we get the relation
@@ Appendix D Systematics from the numerical quadratures
Gauss $22.768(46)$
Gauss-Kronrod $22.756(46)$
@@ E.1 Tuning of the HMC
We performed the tuning of the algorithm parameters at the different values of the hopping parameter $\kappa$, required by the integration in the quark mass, on $6\times 48^{3}$ lattices at one representative value of bare coupling and shift $L_{0}{{\boldsymbol{\xi}}}=(8a,0,0)$. We started from the ensemble at the lowest value of $\kappa$ and we progressively increased it towards the critical value, exploring several algorithmic setups at each value of the hopping parameter. We defined a performance estimator given by the ratio of the computational time over acceptance rate: the best algorithm at given $\kappa$ is the one that minimizes this estimator (with a lower bound of about $90\%$ on the acceptance). We also monitored the spectrum of the Dirac operator so to optimally tune the twisted-masses for the Hasenbusch preconditioning and the RHMC parameters.
@@
At the smallest values of hopping parameter we used a 2-level multistep algorithm where the gauge force is at the finest level and the fermionic forces at the coarsest [71]. The latter include the contributions from the up-down doublet and from the RHMC for the strange quark. Both levels have been integrated with a 4th order Omelyan-Mryglod-Folk (OMF) scheme [72], the finest with 1 step and the coarsest with a tunable number of steps ranging between 7 and 9 for increasing hopping parameter and decreasing temperature. By looking at the performance estimator introduced above and at the acceptance, we observed that the Hasenbusch preconditioning of the up-down doubled determinant is not required up to $\kappa\sim 0.10$. In the interval $0.10\lesssim\kappa\lesssim 0.12$ this simple algorithm is refined with one Hasenbusch twisted-mass equal to $\sqrt{\lambda_{\rm min}\lambda_{\rm max}}$ [73], where $\lambda_{\rm min}$ and $\lambda_{\rm max}$ are respectively the smallest and largest eigenvalues of the operator $\sqrt{D^{\dagger}D}$, whose determination is described later. In this interval the number of steps for the molecular dynamics at the coarsest level ranges between 9 and 15. For $\kappa\gtrsim 0.12$, we opted for a frequency splitting of the RHMC in two terms, and we added a third integration level to profit from the hierarchy in magnitude of these two contributions. The two finest levels have been integrated with 1 step of 4th order OMF scheme, and the coarsest level with a 2nd order OMF scheme with a tunable number of steps from 12 to 17. More elaborated algorithms were not competitive up to $\kappa\gtrsim 0.126$, where we split the up-down doublet determinant with one further Hasenbusch twisted-mass. We tried many setups, including a generalization of the rule proposed in Ref. [73] to the two twisted-mass case. At the end we found that the two values $\lambda_{\rm min}+\frac{1}{3}(\lambda_{\rm max}-\lambda_{\rm min})$ and $\lambda_{\rm max}-\frac{1}{3}(\lambda_{\rm max}-\lambda_{\rm min})$ are a good choice in terms of performance. Here the number of steps for the molecular dynamics at the coarsest level ranges between 13 and 20.
@@ E.2 Generation of ensembles
The lattices with target spatial size $L/a=144$ were generated by tripling the $L/a=48$ ones in the spatial directions, after verifying that the initial smaller-volume configuration was in the zero topological sector and had trivial $\mathbb{Z}_{3}$ phase. At the temperature $T_{8}$, resolution $L_{0}/a=6$ and $\kappa\lesssim 0.10$, we further refined the thermalization procedure considering two steps: from $L/a=24$ lattices to $L/a=72$ ones and finally to the target volume by anti-periodically extending the $L/a=72$ lattices to the $L/a=144$ ones so to start with exactly zero topology at the target volume. The Monte Carlo history of the average plaquette suggested to skip 5050-$100$ trajectories as thermalization, after which we finally measured our observables. The same checks on energy-momentum tensor, topology and $\mathbb{Z}_{3}$ phase have been performed on the large volume ensembles as well.
@@ Appendix G Optimization of the statistics
We discuss here the strategy that we have implemented for the optimization of the computational time needed for the numerical determination of $\frac{\Delta(f_{{\boldsymbol{\xi}}}-f_{{\boldsymbol{\xi}}}^{\infty})}{\Delta\xi_{k}}$}, computed as described in Section IV. At a given set of bare parameters $L_{0}/a$ and $g_{0}^{2}$ the error squared of this contribution is propagated from the fluctuations of the scalar density as follows,

### v2

@@ Abstract
We present a detailed account of the theoretical progress and the computational strategy that led to the non-perturbative determination of the QCD Equation of State at temperatures up to the electroweak scale reported in [Phys. Rev. Lett. 134, 201904 (2025)]. The two key ingredients that make such a calculation feasible with controlled uncertainties are: (i) the definition of lines of constant physics through the running of a non-perturbatively defined finite-volume coupling across a wide range of energy scales, and (ii) the use of shifted boundary conditions which allow a direct determination of the entropy density thus without the need for a zero-temperature subtraction. Considering the case of QCD with $N_{f}=3$ massless flavours in the temperature interval between 3 GeV and 165 GeV, we describe the numerical strategy based on integrating in the bare coupling and quark mass, the perturbative improvement of lattice observables, the optimization of numerical simulations, and the continuum extrapolation. Extensive consistency checks, including finite volume and topology freezing effects, confirm the robustness of the method. The final results have a relative accuracy of about $1\%$ or better, and the errors are dominated by the statistical fluctuations of the Monte Carlo ensembles. We also compare our non-perturbative results with predictions from standard and hard thermal loop perturbation theory showing that at the level of $\%$-precision contributions beyond those known, including non-perturbative ones due to ultrasoft modes, are relevant up to the highest temperatures explored. The methodological framework is general and readily applicable to QCD with four and five massive quark flavours and to other thermal observables, paving the way for systematic non-perturbative studies of thermal QCD at very high temperatures.
@@ I Introduction
The study of the thermodynamics of QCD is central to our understanding of strongly interacting matter under extreme conditions, with direct implications for cosmology [1, 2] and heavy-ion physics [3, 4]. A key challenge is the reliable determination of equilibrium properties from first principles across energy scales where perturbation theory is either ineffective or insufficiently accurate. Lattice QCD has proven to be the most powerful tool in this domain with full control on the statistical and on the systematic effects. However, traditional numerical strategies for computing the Equation of State (EoS) by Monte Carlo simulations face severe limitations when extended to temperatures higher than 1 GeV or so. In particular, ultraviolet subtractions and uncertainties in the setting of the bare parameters to determine the lines of constant physics, complicate the realization of this computation and the extraction of results holding in the continuum limit.
To overcome these obstacles, we have developed novel methodologies. They combine non-perturbative renormalization techniques based on the running of a renormalized finite-volume gauge coupling up to very high energies with shifted boundary conditions in the temporal direction, allowing to directly determinate the entropy density which does not need zero-temperature subtractions. This framework allows precise simulations at high temperatures while keeping the computational cost under control and providing a simple and robust method to extrapolate the results to the continuum limit at a given physical temperature $T$. The approach is general and can be systematically applied to QCD with an arbitrary number of quark flavours, including heavy ones, without further theoretical adjustments.
@@ V.1 Finite volume effects
are defined as in Section 3 of Ref. [10]. At present, the entropy density cannot be determined non-perturbatively from Eq. (16) because the renormalization constants $Z_{G}^{\{6\}}$ and $Z_{F}^{\{6\}}$ are known only at one-loop order in bare lattice perturbation theory. However, since the latter are insensitive to finite volume effects, the matrix elements ${\langle T_{\mu\nu}^{G,\{6\}}\rangle}_{{\boldsymbol{\xi}}}$ and ${\langle T_{\mu\nu}^{F,\{6\}}\rangle}_{{\boldsymbol{\xi}}}$ can be used as a proxy to study the finite volume effects affecting the entropy density. Their lattice determination is significantly less demanding than the computation of the entropy via Eq. (6).
@@ VI Equation of State
The main non-perturbative results are the values of the normalized entropy density $s/T^{3}$ of continuum QCD reported in Table 5 with the fit label id3, corresponding to the 9 temperatures from 3 GeV to 165 GeV of Table 1. These values are shown as black dots in Figure 5 as a function of the strong coupling squared $\hat{g}^{2}$ defined above. These results can be fitted to obtain an analytic expression for the temperature dependence of the entropy density, and thus for the other thermodynamic functions in the EoS, namely the pressure and the energy density.
@@ VI.1 Temperature dependence of the entropy density
In the second fit we have enforced all the known coefficients in perturbation theory up to $O(\hat{g}^{6}\ln\hat{g}^{2})$  [38, 39, 40, 41, 42, 43, 44, 45] so as to constrain the parametrization in the asymptotically high temperature regime. We have fitted to the data the unknown term at $O(\hat{g}^{6})$ and one further term of $O(\hat{g}^{7})$ to take into account any higher order effects in the expansion in $\hat{g}$. The green band in Figure 5 shows this fit. The fitted coefficients actually include the non-perturbative contribution of the ultrasoft modes [46, 47] and a not yet known perturbative term. The determination of the missing perturbative coefficient at $O(\hat{g}^{6})$ (see Ref. [48] for recent progresses) would allow us to disentangle the ultrasoft effects and estimate the relative magnitude.
@@
Figure 7: Left: Normalized entropy density $s/T^{3}$ in the continuum limit as a function of $\hat{g}^{2}$. The blue band is the best parametrization for $T\geq 500$ MeV, and includes the point at $T=500$ MeV obtained from Refs. [49, 50, 32] (open circle marker). Right: EoS as a function of the temperature for $T\geq 500$ MeV.
@@ VII Conclusions
A detailed analysis of discretization effects, of finite volume corrections and of topological contributions confirms the robustness of our results. We have also performed a comparison with perturbative predictions, including both standard and hard thermal loop resummations. Our findings show that, at our level of accuracy, higher-order contributions in $\hat{g}^{2}$ beyond those known, including non-perturbative ones due to ultrasoft modes, are relevant even at temperatures as high as 165 GeV. Furthermore, our results are consistent with existing lattice data at lower temperatures, down to 500 MeV, and a smooth interpolation can be drawn across the entire range.
The methodology discussed in this paper is broadly applicable and can be extended to QCD with four or five massive quark flavours. By disentangling the methodological framework from its first physical application, this work complements earlier results and sets the stage for a systematic program of non-perturbative studies of QCD in the high-temperature regime, with potential impact on cosmology, heavy-ion phenomenology, and the theoretical development of thermal field theory. In particular, the knowledge of the derivative of the free-energy in the shift paves the way to the non-perturbative definition of the QCD energy-momentum tensor, following the strategy proposed in Ref. [10]. Solving this long standing theoretical problem [56, 57, 58, 59] would enable further first-principles investigations of the thermal properties of QCD through correlation functions of the energy-momentum tensor.
@@ Appendix B Free-energy density in lattice perturbation theory
In this Appendix we discuss the results of the free-energy density at one-loop order in lattice perturbation theory, evaluated in the infinite spatial volume limit. These expressions, taken from Ref. [10] (in particular Appendices E and F), are used for the perturbative improvement of the entropy density computed non-perturbatively on the lattice.
@@ B.3.2 Integrals at one-loop order
In left panel of Figure 8 we compare the approach to the thermodynamic limit of $K_{1}^{\prime}$ (orange square markers) and $K_{1}$ (blue circle markers). The system size in the spatial directions increases from $L/a=96$ up to $L/a=288$. For all practical purposes we can quote as the infinite volume limit the value of $K_{1}^{\prime}$ at the largest volume, whose discrepancy with respect to the extrapolation is well below the permille level.
The leading finite size corrections of the integral $K_{2}$ are of order $O(1/L^{3})$ and they are negligible at the largest size $L/a=288$ that we have considered. Thus, we take that value as the thermodynamic limit.
@@ B.4 Infinite volume limit of fermionic contributions
We have extrapolated to the thermodynamic limit the tree-level fermionic free-energy density $f^{F(0)}$, Eq. (58), and the one-loop order mass counterterm, Eq. (67), by increasing $L/a$ from $96$ to $288$ at fixed $L_{0}/a$. The convergence to the infinite volume limit is very fast because, at asymptotically high temperatures, fermions develop a thermal mass $M_{\rm th}\sim\pi T$ and finite volume effects are exponentially suppressed as $e^{-M_{\rm th}L}$ [9]. At the largest volume, finite size effects are well below the permille level, and in practice the results can be considered as the values in the thermodynamic limit.
@@
whose integrand has a pole when the gluonic momentum vanishes $q\to 0$ and $p\sim k$. Following the same strategy as for the computation of $K_{1}$ in the previous Section, we have introduced the subtracted integral
@@ C.2 Relative error
In the presence of shifted boundary conditions, the temperature of the thermal system depends on the shift through the relation $T^{-1}=L_{0}\sqrt{1+{{\boldsymbol{\xi}}}^{2}}$. At a given $L_{0}$, the temperature changes by less than a factor 2 with respect to the reference case ${{\boldsymbol{\xi}}}=(1,0,0)$ when the shift is varied within the values listed in Eq. (78). Since the entropy density increases slowly from 3 GeV to 165 GeV (about $3\%$, see Table 5), at these temperatures and at a given $L_{0}$ we can assume $s/T^{3}$ to be approximately constant when the shift is changed among the values under investigation. Using this assumption in Eq. (77), we get the scaling
@@ Appendix D Systematics from the numerical quadratures
Gauss $22.77(5)$
Gauss-Kronrod $22.76(5)$
@@ E.1 Tuning of the HMC
We performed the tuning of the algorithm parameters at the different values of the hopping parameter $\kappa$, required by the integration in the quark mass, on $6\times 48^{3}$ lattices at one representative value of bare coupling and shift $L_{0}{{\boldsymbol{\xi}}}=(8a,0,0)$. We started from the ensemble at the lowest value of $\kappa$ and we progressively increased it towards the critical value, exploring several algorithmic setups at each value of the hopping parameter. We defined a performance estimator given by the ratio of the computational time over acceptance rate: the best algorithm at given $\kappa$ is the one that minimizes this estimator (with a lower bound of about $90\%$ on the acceptance). We also monitored the spectrum of the Dirac operator so as to optimally tune the twisted-masses for the Hasenbusch preconditioning and the RHMC parameters.
@@
At the smallest values of hopping parameter we used a 2-level multistep algorithm where the gauge force is at the finest level and the fermionic forces are at the coarsest one [74]. The latter include the contributions from the up-down doublet and from the RHMC for the strange quark. Both levels have been integrated with a 4th order Omelyan-Mryglod-Folk (OMF) scheme [75], the finest with 1 step and the coarsest with a tunable number of steps ranging between 7 and 9 for increasing hopping parameter and decreasing temperature. By looking at the performance estimator introduced above and at the acceptance, we observed that the Hasenbusch preconditioning of the up-down doublet determinant is not required up to $\kappa\sim 0.10$. In the interval $0.10\lesssim\kappa\lesssim 0.12$ this simple algorithm is refined with one Hasenbusch twisted-mass equal to $\sqrt{\lambda_{\rm min}\lambda_{\rm max}}$ [76], where $\lambda_{\rm min}$ and $\lambda_{\rm max}$ are respectively the smallest and largest eigenvalues of the operator $\sqrt{D^{\dagger}D}$, whose determination is described later. In this interval the number of steps for the molecular dynamics at the coarsest level ranges between 9 and 15. For $\kappa\gtrsim 0.12$, we opted for a frequency splitting of the RHMC in two terms, and we added a third integration level to profit from the hierarchy in magnitude of these two contributions. The two finest levels have been integrated with 1 step of 4th order OMF scheme, and the coarsest level with a 2nd order OMF scheme with a tunable number of steps from 12 to 17. More elaborated algorithms were not competitive up to $\kappa\gtrsim 0.126$, where we split the up-down doublet determinant with one further Hasenbusch twisted-mass. We tried many setups, including a generalization of the rule proposed in Ref. [76] to the two twisted-mass case. At the end we found that the two values $\lambda_{\rm min}+\frac{1}{3}(\lambda_{\rm max}-\lambda_{\rm min})$ and $\lambda_{\rm max}-\frac{1}{3}(\lambda_{\rm max}-\lambda_{\rm min})$ are a good choice in terms of performance. Here the number of steps for the molecular dynamics at the coarsest level ranges between 13 and 20.
@@ E.2 Generation of ensembles
The lattices with target spatial size $L/a=144$ were generated by tripling the $L/a=48$ ones in the spatial directions, after verifying that the initial smaller-volume configuration was in the zero topological sector and had trivial $\mathbb{Z}_{3}$ phase. At the temperature $T_{8}$, resolution $L_{0}/a=6$ and $\kappa\lesssim 0.10$, we further refined the thermalization procedure considering two steps: from $L/a=24$ lattices to $L/a=72$ ones and finally to the target volume by anti-periodically extending the $L/a=72$ lattices to the $L/a=144$ ones so as to start with exactly zero topology at the target volume. The Monte Carlo history of the average plaquette suggested to skip 5050-$100$ trajectories as thermalization, after which we finally measured our observables. The same checks on energy-momentum tensor, topology and $\mathbb{Z}_{3}$ phase have been performed on the large volume ensembles as well.
@@ Appendix G Optimization of the statistics
We discuss here the strategy that we have implemented for the optimization of the statistics needed for the numerical determination of $\frac{\Delta(f_{{\boldsymbol{\xi}}}-f_{{\boldsymbol{\xi}}}^{\infty})}{\Delta\xi_{k}}$}, computed as described in Section IV. At a given set of bare parameters $L_{0}/a$ and $g_{0}^{2}$ the error squared of this contribution is propagated from the fluctuations of the scalar density as follows,